If x and y are positive consecutive odd integers, where y > x. which of the following is equal to y^2-x^2?
(A)2x
(B)4x
(c)2x+2
(D)2x+4
(E)4x+4

Answer by ankor@dixie-net.com(22740)

If x and y are positive consecutive odd integers, where y > x. which of the following is equal to y^2-x^2?
Then y = (x+2)
:
Substitute for y
(x+2)^2 - x^2
FOIL
x^2 + 4x + 4 - x^2
x^2 - x^2 + 4x + 4
then
4x + 4 = y^2 - x^2
: